A transverse wave is describe by the function y=(.08 m) sin[(28 radians) t+(5.6 m^-1)x]. Find the amplitude, angular frequency, frequency, period, wavelength and speed of the wave.
1
Expert's answer
2021-02-24T12:48:55-0500

a)

"A=0.08\ m."

b)

"\omega=28\ \dfrac{rad}{s}."

c)

"f=\dfrac{\omega}{2\pi}=\dfrac{28\ \dfrac{rad}{s}}{2\pi}=4.46\ Hz."

d)

"T=\dfrac{1}{f}=\dfrac{1}{4.46\ Hz}=0.224\ s."

e)

"\lambda=\dfrac{2\pi}{k}=\dfrac{2\pi}{5.6\ m^{-1}}=1.12\ m."

f)

"v=f\lambda=4.46\ Hz\cdot1.12\ m=5\ \dfrac{m}{s}."